# Phase Transitions in Anisotropic Turbulence

### v1

@@ IV.3 Transitions between distinct large-scale flows
Figure 6: Stochasic, rare transitions between ‘hurricane-like’ large-scale vortices and unidirectional jets (not unlike the Earth’s jet stream) in 2D turbulence within a weakly elongated, periodic domain (at $\Delta=1.06$, i.e. $6\%$ elongation in the $x$-direction). Bottom: snapshots of $\omega_{z}$ in large-scale vortex (left) and jet (right) states.
@@ V.1 Turbulence over topography
Quasi-2D flows above topography are both a standard model for various large-scale oceanic flows and provide a testbed for studying how spatial disorder disrupts the large-scale self organization of two-dimensional flows. Interesting oceanic examples include the Antarctic Circumpolar Current, where turbulent eddies are channeled by and may be locked to the bottom topography Straub (1993) and, at smaller scales, the Lofoten basin vortex,Søiland and Rossby (2013) a highly persistent anticyclone located at the center of the basin.
@@
Early on, the study of quasi-2D flows over topography was interlinked with the search for general principles governing large-scale self-organization in quasi-2D flows more generally. Bretherton and HaidvogelBretherton and Haidvogel (1976) applied the concept of ‘selective decay’ (an organizational principle borrowed from plasma physics) to such flows. The selective decay principle states that forward-cascading invariants should be minimized for a given (conserved) value of the inverse-cascading invariants. Based on the inverse energy cascade and forward enstrophy cascade of 2D turbulence, Bretherton and Haidvogel thus conjectured that the system evolves towards a state of minimum enstrophy, while conserving its initial energy. The limitation of this and similar results, as in homogeneous 2D turbulence, is that they are formulated based on exact energy conservation, and may not apply in the presence of forcing or dissipation.
@@
Recently, Siegelman and YoungSiegelman and Young (2023) revisited the problem of unforced and weakly damped quasi-geostrophic flow over random rough topography by means of numerical simulations. In qualitative agreement with selective decay, enstrophy decreased with time in their simulations, while energy was conserved to a good approximation. However, the flows Siegelman and Young observed in the long-time limit depart from the predictions of selective decay. Specifically, they identified striking differences between flows with low kinetic energy $\mathcal{E}$ and flows with high $\mathcal{E}$, separated by a threshold energy $\mathcal{E}_{\rm t}$ associated with a given topography: flows with $\mathcal{E}<\mathcal{E}_{\rm t}$ feature (weak) vortices which are strongly correlated with the bottom topography, with anticyclones locked to topographic depressions and cyclones to topographic elevations, a phenomenon not predicted by selective decay. At $\mathcal{E}>\mathcal{E}_{\rm t}$, strong vortices are able to detach themselves from the topography and the vorticity-topography correlation disappears.
@@
In a follow-up study, GalletGallet (2024) investigated the case where topography is present only at scales smaller than the size of the domain, unlike Bretherton and Haidvogel or Siegelman and Young, who had assumed topography at all scales. Gallet showed that there is in fact a phase transition at the critical energy $\mathcal{E}=\mathcal{E}_{\rm t}$ between a state without large-scale structure at low $\mathcal{E}$ and an inverse cascade with associated large-scale condensation at high $\mathcal{E}$ in this system (cf. Fig 7), generalizing the selective decay principle to include the large-scale condensate, explaining his findings. Very recent resultsZhang and Xie (2024) on quasi-geostrophic flow over small-scale topography revealed even richer behavior than that uncovered by Gallet, with two threshold values of the topography amplitude and bidirectional energy transfers.
(a)         (b)         (c)
@@
Figure 7: Panel (a): example of single-scale topography over which 2D turbulence was studied in.Gallet (2024) Panel (b): snapshot of streamfunction for low-energy initial conditions ($\mathcal{E}<\mathcal{E}_{\rm t}$), showing a strong correlation with the topography. Panel (c): same as (b) for a high-energy initial condition ($\mathcal{E}>\mathcal{E}_{t})$, where the correlation with the topography is absent and a large-scale condensate has formed. Visualizations are adapted from Gallet (2024).
@@ V.2 Instability-driven two-dimensional turbulence
To investigate the impact of the forcing-flow correlation typical of an instability-type forcing mechanism (but absent for stochastic stirring), following earlier work,Jiménez and Guegan (2007) van Kan et al.van Kan et al. (2022, 2024) recently studied 2D turbulence driven by a superposition of two forces $\mathbf{f}=\mathbf{f}_{\gamma}\equiv(1-\gamma)\mathbf{f}_{\rm stoch}+\gamma\mathbf{f}_{\rm inst}$, parameterized by a control parameter $\gamma\in[0,1]$. The first term is a stochastic force injecting a constant power within a fixed (thin, small-scale) band consisting of wave numbers $[k_{1},k_{2}]$, independently of the flow state, while the second term is a linear operator acting on the velocity field which is filtered to retain only the same wave numbers $[k_{1},k_{2}]$ as the stochastic term. To saturate the linear instability resulting from the forcing, a nonlinear damping was included in addition to hyperviscosity.
@@
For small values of $\gamma$, where the random forcing dominates, integrating the equations from rest yields an inverse energy cascade producing a large-scale condensate (Fig. 8(a)). By contrast, as $\gamma$ is increased (stronger instability forcing) the flow morphology changes drastically, and coherent vortices emerge, which are tripolar (consisting of a vorticity extremum in the core and two satellites of the opposite sign at 180 apart). These coherent vortices are also shielded in the sense that the circulation they induce is close to zero outside a finite radius (slightly larger than the forcing scale). At intermediate values of $\gamma$, shielded vortices of both signs persist, embedded inside the large-scale vortices (Fig. 8(b)). At larger $\gamma$, another phase transition occurs, where the condensate is suppressed and the vortices undergo a spontaneous symmetry breaking, with all vortices of one (arbitrary) sign being annihilated, leading to a polarized vortex gas (Fig. 8(c)). The density of this vortex gas slowly increases in time due to random nucleation, up to a density threshold where an explosive nucleation phase leads to a final state consisting of a high-density gas. This dense vortex gas can in turn be continued to smaller $\gamma$ (weaker instability forcing), where, at a critical forcing strength $\gamma=\gamma_{c}$, it forms a hexagonal vortex crystal (Fig. 8(d)), not unlike those found in rotating flows,Boubnov and Golitsyn (1986); Di Leoni et al. (2020) at the Jovian poles,Adriani et al. (2018) and in active matter.Riedel, Kruse, and Howard (2005); Xu and Wu (2024) A suitable order parameter for the crystallization transition is the diffusivity of individual shielded vortices, which undergo a random Brownian-like motion in the gas but are trapped in the crystal. The diffusivity shows a power law dependence on $\gamma-\gamma_{c}$, which has not yet been explained theoretically.
@@
These results highlight the forcing dependence of 2D turbulence, implying non-universality.Linkmann, Hohmann, and Eckhardt (2020) It is anticipated that the properties of instability-driven turbulence described above can be verified in more complex quasi-2D flows driven by spectrally localized instabilities beyond this simple model. However, many questions remain open, regarding the robustness of the observed phases to changes in control parameters other than $\gamma$, as well as the stability of the shielded vortices in a quasi-2D configuration, and the theoretical description of the different transitions (including crystallization) observed in this forced-dissipative system. Moreover, while many features of inertial instability-driven turbulence at high $\rm Re$ resemble (low-$\rm Re$) active fluid flows, including at the level of the mathematical models and the observed flow states, the precise relation between these two physically distinct problems remains to be clarified.
@@
Figure 8: Illustrations (showing $\omega_{z}$) of different flow states observed in instability-driven 2D turbulence. (a): large-scale codnensate, (b) hybrid state with large-scale condensate and embedded shielded vortices of either sign. (c): shielded vortex gas of broken symmetry. (d): vortex crystal. Figure adapted from van Kan et al. (2024)van Kan et al. (2024).
@@ VI Conclusions
In this mini-review, we have surveyed recent developments in the area of phase transitions in highly anisotropic turbulent flows. We have described how turbulence transitions from 3D to quasi-2D and 2D in different systems, and discussed key examples of multistability and turbulent noise-induced transitions. Finally, we reviewed recently discovered phase transitions in modified forms of 2D turbulence.
The problems described above are of course highly idealized and pale in comparison to the full complexity of the Earth System. However, they nonetheless pose significant challenges to our understanding and allow us to learn about the fundamental processes of turbulence which govern not only the transport of heat, momentum, but also pollutants and nutrients on our planet and beyond. Going forward, the insights gained about geophysical and astrophysical turbulence in the process may help inform parameterizations of turbulence in weather and climate models, which do not typically take inverse energy cascades into account.
Finally, it is interesting to note that the phase transitions discussed here can be viewed in some sense as a form of tipping pointsLenton (2011) in turbulence, namely, thresholds in parameter space, which, when crossed, induce a rapid, qualitative change in the global system state.
@@ Acknowledgements
This work was supported by the National Science Foundation (grants DMS-2009563 and DMS-2308337) and the German Research Foundation (DFG Projektnummer: 522026592). Part of the writing of this article was done while the author was a Staff Member at the Woods Hole Oceanographic Institution Summer Program in Geophysical Fluid Dynamics summer program 2024, with partial support from NSF (grant DMS-2308338). The author thanks Alexandros Alexakis, Santiago Benavides, Lichuan Xu and Edgar Knobloch for their feedback on earlier versions of this manuscript and Chia-Chin Chung for her support with graphic design.

### v2

@@ IV.3 Transitions between distinct large-scale flows
[…]
@@ V.1 Turbulence over topography
The following section is largely inspired by the work of GalletGallet (2024), with the permission of the author. Highly anisotropic fluid flows above topography provide an idealized model for large-scale geophysical flows and also represent a proving ground for studying how spatial inhomogeneity impacts large-scale structure formation in (quasi-)two-dimensional flows. Interesting oceanic examples include the Antarctic Circumpolar Current, where turbulent eddies are funneled by and may be locked to bottom topography Straub (1993) and, at smaller scales, the Lofoten basin vortex,Søiland and Rossby (2013) a highly persistent anticyclone located at the center of the basin.
@@
The study of quasi-geostrophic flows, in which the horizontal momentum balance is dominated by the Coriolis force associated with planetary rotation balanced by pressure gradients and advection is primarily horizontal, over bottom topography has long been was interlinked with the broader study of large-scale self-organization in highly anisotropic flows. Bretherton and HaidvogelBretherton and Haidvogel (1976) applied the concept of ‘selective decay’ (an organizational principle borrowed from plasma physics) to such flows, which states that forward-cascading invariants should be minimized for a given value of the inverse-cascading invariants. Based on the inverse energy cascade and forward enstrophy cascade of 2D turbulence (both of which also arise in quasi-geostrophic turbulence), Bretherton and Haidvogel thus predicted that the system evolves towards a state characterized by a minimum in the enstrophy, while conserving its initial energy. The limitation of this and similar results, as in homogeneous 2D turbulence, is that they are formulated based on exact energy conservation, and may not apply in the presence of forcing or dissipation.
@@
Siegelman and YoungSiegelman and Young (2023) recently reconsidered the problem of unforced and weakly damped quasi-geostrophic flow over random rough topography by means of numerical simulations. As one would expect from selective decay, enstrophy decreased over time in their simulations, while energy was approximately constant. However, at late times, the flows observed by Siegelman and Young departed from the predictions of selective decay. Specifically, they identified striking differences between flows with low kinetic energy $\mathcal{E}$ and flows with high $\mathcal{E}$, separated by a threshold energy $\mathcal{E}_{\rm t}$ associated with a given topography: flows with $\mathcal{E}<\mathcal{E}_{\rm t}$ feature (weak) vortices which are strongly correlated with the bottom topography, with anticyclones locked to topographic depressions and cyclones to topographic elevations, a phenomenon not predicted by selective decay. At $\mathcal{E}>\mathcal{E}_{\rm t}$, strong vortices are able to detach themselves from the topography and the vorticity-topography correlation disappears.
@@
In a follow-up study, GalletGallet (2024) investigated the case where topography is present only at scales smaller than the size of the domain, unlike Bretherton and Haidvogel or Siegelman and Young, who had assumed topography at all scales. Gallet showed that there is in fact a phase transition at the critical energy $\mathcal{E}=\mathcal{E}_{\rm t}$ between a state without large-scale structure at low $\mathcal{E}$ and an inverse cascade with associated large-scale condensation at high $\mathcal{E}$ in this system (cf. Fig. 9), generalizing the selective decay principle to include the large-scale condensate, explaining his findings. Very recent resultsZhang and Xie (2024) on quasi-geostrophic flow over small-scale topography revealed even richer behavior than that uncovered by Gallet, with two threshold values of the topography amplitude and bidirectional energy transfers.
(b)         (c)         (d)
@@
Figure 9: Panel (a): Nondimensional condensate amplitude versus total energy revealing phase transition in large-scale structure, in agreement with generalized selective decay, with only a weak dependence on initial conditions. Panel (b): example of single-scale topography over which 2D turbulence was studied in.Gallet (2024) Panel (c): snapshot of streamfunction for low-energy initial conditions ($\mathcal{E}<\mathcal{E}_{\rm t}$), showing a strong correlation with the topography. Panel (d): same as (c) for a high-energy initial condition ($\mathcal{E}>\mathcal{E}_{t})$, where the correlation with the topography is absent and a large-scale condensate has formed. Visualizations are adapted from Gallet (2024). Source: Basile Gallet, J. Fluid Mech., 988, A13, 2024; licensed under a Creative Commons Attribution (CC BY) license.
@@ V.2 Instability-driven two-dimensional turbulence
To investigate the impact of the forcing-flow correlation typical of an instability-type forcing mechanism (but absent for stochastic stirring), following earlier work,Jiménez and Guegan (2007) van Kan et al.van Kan et al. (2022, 2024b) recently studied 2D turbulence driven by a superposition of two forces $\mathbf{f}=\mathbf{f}_{\gamma}\equiv(1-\gamma)\mathbf{f}_{\rm stoch}+\gamma\mathbf{f}_{\rm inst}$, parameterized by a control parameter $\gamma\in[0,1]$. The first term is a stochastic force injecting a constant power within a fixed (thin, small-scale) band consisting of wave numbers $[k_{1},k_{2}]$, independently of the flow state, while the second term is a linear operator acting on the velocity field which is filtered to retain only the same wave numbers $[k_{1},k_{2}]$ as the stochastic term. To saturate the linear instability resulting from the forcing, a nonlinear damping was included in addition to hyperviscosity.
@@
For small values of $\gamma$, where the random forcing dominates, integrating the equations from rest yields an inverse energy cascade producing a large-scale condensate (Fig. 10(a)). By contrast, as $\gamma$ is increased (stronger instability forcing) the flow morphology changes drastically, and coherent vortices emerge, which are tripolar (consisting of a vorticity extremum in the core and two satellites of the opposite sign at 180 apart). These coherent vortices are also shielded in the sense that the circulation they induce is close to zero outside a finite radius (slightly larger than the forcing scale). At intermediate values of $\gamma$, shielded vortices of both signs persist, embedded inside the large-scale vortices (Fig. 10(b)). At larger $\gamma$, another phase transition occurs, where the condensate is suppressed and the vortices undergo a spontaneous symmetry breaking, with all vortices of one (arbitrary) sign being annihilated, leading to a polarized vortex gas (Fig. 10(c)). The density of this vortex gas slowly increases in time due to random nucleation, up to a density threshold where an explosive nucleation phase leads to a final state consisting of a high-density gas. This dense vortex gas can in turn be continued to smaller $\gamma$ (weaker instability forcing), where, at a critical forcing strength $\gamma=\gamma_{c}$, it forms a hexagonal vortex crystal (Fig. 10(d)), not unlike those found in rotating flows,Boubnov and Golitsyn (1986); Di Leoni et al. (2020) at the Jovian poles,Adriani et al. (2018) and in active matter, both in experimentsRiedel, Kruse, and Howard (2005); Xu and Wu (2024) and simulations.James, Bos, and Wilczek (2018); James et al. (2021). A suitable order parameter for the crystallization transition is the diffusivity of individual shielded vortices, which undergo random Brownian-like motion in the gas phase but are trapped in the crystal phase. The diffusivity shows a power law dependence on $\gamma-\gamma_{c}$, which has not yet been explained theoretically.
@@
These results highlight the forcing dependence of 2D turbulence, implying non-universality.Linkmann, Hohmann, and Eckhardt (2020) It is anticipated that the properties of instability-driven turbulence described above can be verified in more complex quasi-2D flows driven by spectrally localized instabilities beyond this simple model. However, many questions remain open, regarding the robustness of the observed phases to changes in control parameters other than $\gamma$, as well as the stability of the shielded vortices in a quasi-2D configuration, and the theoretical description of the different transitions (including crystallization) observed in this forced-dissipative system. Moreover, while many features of inertial instability-driven turbulence at high $\rm Re$ resemble (low-$\rm Re$) active fluid flows, including at the level of the mathematical models (Swift-Hohenberg-Toner-Tu equationsAlert, Casademunt, and Joanny (2022)) and the observed flow states, the precise relation between these two physically distinct problems remains to be clarified.
@@
Figure 10: Illustrations (showing $\omega_{z}$) of different flow states observed in instability-driven 2D turbulence. (a): large-scale condensate, (b) hybrid state with large-scale condensate and embedded shielded vortices of either sign. (c): shielded vortex gas of broken symmetry. (d): vortex crystal. Reproduced with permission from J. Fluid Mech. 380:20210049 (2024). Copyright 2024 Cambridge University Press.
@@ VI Conclusions
In this mini-review, we have surveyed recent developments in the area of phase transitions in highly anisotropic turbulent flows. We have described how turbulence transitions from 3D to quasi-2D and 2D in different systems. In the presence of large-scale damping, the transition towards the inverse energy cascade is continuous (i.e., second-order) in terms of the order parameter given by the inverse energy flux. The scaling of the inverse energy flux with the physical control parameter near onset of the inverse cascade has been found to be approximately linear in several cases including thin-layer turbulence, but more data is needed and no theory exists to predict the behavior of the system in this parameter regime. When large-scale damping is absent, the transition from small-scale 3D turbulence to the large-scale condensate is discontinuous (i.e., first-order) in terms of the large-scale kinetic energy in the thin-layer and rotating turbulence cases. If the anisotropy is increased sufficiently, e.g., by reducing the layer depth to the Kolmogorov scale, the kinetic energy in 3D modes vanishes with a different power law, whose exponent appears to be larger than one, which also remains incompletely understood. Importantly, nearly all studies summarized above adopted periodic boundary conditions in the horizontal and vertical directions. How these transitions are impacted by no-slip boundary conditions, as they are encountered in laboratory experiments, remains yet to be studied. The addition of the competing effects of rotation and stratification enriches the dynamics near the onset of the inverse cascade and more work remains to be done to characterize the large-scale vortices emerging in this case. We also discussed key examples of multistability and turbulent noise-induced transitions and reviewed recently discovered phase transitions in modified forms of 2D turbulence, where approaches from equilibrium statistical theory have proven successful in some cases, but the limitations of such methods to forced-dissipative turbulence more generally remain incompletely understood.
The problems described above are of course highly idealized and pale in comparison to the full complexity of the Earth System. However, they nonetheless pose significant challenges to our understanding and allow us to learn about the fundamental processes of turbulence, including the direction and rate of energy transfers across scales and the associated formation or absence of large-scale flow structures, which govern not only the transport of heat, momentum, but also pollutants and nutrients on our planet and beyond. Going forward, the insights gained about geophysical and astrophysical turbulence in the process may help inform parameterizations of turbulence in weather and climate models, which do not typically take inverse energy cascades into account.
Finally, it is interesting to note that the phase transitions discussed here can be viewed as a form of tipping pointsLenton (2011) in turbulence, namely, thresholds in parameter space, which, when crossed, induce a rapid, qualitative change in the global system state. As the control parameters (e.g., of the Earth system) changes slowly, the response in the resulting turbulent fluid flows may be sudden.
@@ Acknowledgements
The author thanks two anonymous referees for their highly constructive reviews which have significantly improved this manuscript. This work was supported by the National Science Foundation (grants DMS-2009563 and DMS-2308337) and the German Research Foundation (DFG Projektnummer: 522026592). Part of the writing of this article was done while the author was a Staff Member at the Woods Hole Oceanographic Institution Summer Program in Geophysical Fluid Dynamics summer program 2024, with partial support from NSF (grant DMS-2308338). The author thanks Alexandros Alexakis, Santiago Benavides, Lichuan Xu and Edgar Knobloch for their feedback on earlier versions of this manuscript and Chia-Chin Chung for her support with graphic design.